A Space & astronomy forum. SpaceBanter.com

Go Back   Home » SpaceBanter.com forum » Space Science » Space Station
Site Map Home Authors List Search Today's Posts Mark Forums Read Web Partners

ISS On-Orbit Status, 19-03-2004



 
 
Thread Tools Display Modes
Prev Previous Post   Next Post Next
  #1  
Old March 19th 04, 08:58 PM
Jacques van Oene
external usenet poster
 
Posts: n/a
Default ISS On-Orbit Status, 19-03-2004

ISS On-Orbit Status 19 Mar 2004

ISS systems continue to function nominally, except those noted previously
or below.

Update on Elektron: After replacing the Liquid Unit #6 (BZh-6) with the
remaining spare (BZh-7) and reconnecting BITS2-12 telemetry cables
yesterday, the crew today replaced feed hoses and cleaned the assembly.
Elektron activation will be attempted tomorrow morning (~2:30am EST).
[Both the Elektron's air/liquid separator unit (GZhS) and the removed parts
(BP pump unit, A-R water feed hoses) were flushed with water from Rodnik
storage tank 1 (BV1), toweled clean and inspected for possible reuse or else
disposal. The flush water was tested for pH (reported value: ~8). The EDV
container with the unclean flush water will be discarded.]

Update on SFOG (TGK) candles: As of tonight, a total of 13 candles have
been burnt since 3/13, all of them without failure. There are 108 candles
left on board, for ~54 days, assuming 100% functionality. The on-orbit
candles are the only ones remaining of the current design. New cartridges,
updated for improved safety (e.g., electrical ignition instead of percussion
cap), are scheduled to enter service late this year (on Progress 16P, unless
made available earlier). The replacement of the fan with the failed bearing
in slot 2 of the TGK combustion container will be performed ASAP. Until
then, candles can be burned only sequentially, not in parallel, with a
3-hour waiting period between ops to allow cool-down of the used SFOG
cassette.

Update on Lab science window: A window leak check today indicated that the
"Volume D" between the window's pressure panes, which was evacuated on 3/5
after the U-jumper flexhose removal, is not holding its vacuum, i.e. is not
airtight (further details see below).

FE Alexander Kaleri cleaned the three IDZh-2 smoke detectors in the DC-1
docking compartment (installed on 9/14/03). [Since the DC-1 does not yet
have the capability to inhibit the IDZh-2's fire response, TsUP/Moscow
partially deactivated the DC-1 onboard computer system (BVS) that normally
monitors the serviceability of each detector. Afterwards, the BVS was
reactivated. The smoke detectors work with electro-induction, i.e., by
first ionizing air molecules in the detector chamber with radioactive
material, then measuring their electric conductivity (which would be reduced
if the air contains smoke particles).]

CDR/SO Michael Foale worked on the MSG PFMI (Microgravity Science
Glovebox/Pore Formation & Mobility Investigation) experiment which had
completed its new sample, PFMI-02, after POIC (Payload Operations &
Integration Center) altered procedures upon the booster heater's not
properly heating up. [Mike removed PFMI-02 from the thermal chamber,
exchanged the PFMI g-LIMIT bypass micro drive assembly in the laptop and
removed the video tape. The MSG was then powered down. The SO also checked
cable matings that may have caused the hitch.]

In the Service Module (SM), Alex Kaleri performed a checkout of the Russian
LIV video system (TVS), in particular evaluating the picture quality of its
two cameras (V/K1, V/K2) and reconfiguring camera and microphone
connections. [Because of degraded picture performance of V/K2, Alex
replaced it with V/K1, after evaluating the latter's picture quality,
recording imagery on the LIV video recorder on a videocassette for return to
the ground.]

The crew conducted their first ultrasound scanning activity with the ADUM
(Advanced Diagnostic Ultrasound in Microgravity) experiment, with Sasha
Kaleri acting as the subject for the experiment and Mike setting up and
using the scanner. [The Ultrasound Imaging System of the HRF (Human
Research Facility) provides three-dimensional image enlargement of the heart
and other organs, muscles and blood vessels. It is capable of high
resolution imaging in a wide range of applications, both research and
diagnostic, such as echocardiography (ultrasound of the heart), abdominal
and deep organ ultrasound, vascular ultrasound, gynecological ultrasound,
muscle and tendon ultrasound, transcranial ultrasound, ultrasound contrast
studies, and small parts ultrasound. Voice & video transmissions involving
human subjects are restricted to honor data confidentiality agreements.
Ultrasound investigations on ISS were done as early as 2001 (Expedition
2/Yuri Usachev, Jim Voss, Susan Helms).]

Foale performed a leak check on the recently depressurized Lab's science
window, with the purpose to verify seal integrity after the inter-pane
volume evacuation on 3/5 and to confirm the actual leak rate. [Equipment
used include FSS (fluid system servicer) and MPEV (manual pressure
equalization valve) adaptors, the Aeolus scopemeter, a timer/stopwatch and
earplugs, if desired. Pressure was measured at 359 mmHg (Torr). After
completion of the evacuation two weeks ago, the pressure had been 2.1 mmHg.
A failure investigation team (FIT) will discuss the results and any
troubleshooting activities on Monday.]

The FE performed his regular maintenance on the BIO-5 Rasteniya-2
("Plants-2") greenhouse. Replenishment of its water supply as required has
now become part of the Russian task list. [Rasteniya studies growth and
development of plants (peas) under spaceflight conditions in the Lada-4
greenhouse. Regular maintenance involves monitoring of seedling growth,
humidity measurements, watering to moisten the substrate if necessary, and
photo/video recording.]

The CDR took the periodic reading of the cabin air's current CO2 partial
pressure in the SM and Lab, using the U.S. CDMK (CO2 monitor kit), for
calldown to MCC-Houston (along with the battery status) for use in trending
analyses.

In the SM, Kaleri corrected misleading labeling on valves of STR condensate
removal pipes, bringing them into line with ground documentation.

The Science Officer did the visual T+2d analysis of Week 21 water samples,
which he collected on 3/17 from the SRV-K hot tap and the EDV container of
the water supply system (SVO-ZV) with the WMK (water monitoring kit). He
then entered the microbiological data in the MEC (medical equipment
computer). [The T+2 analysis uses incubated MCDs (microbial capture
devices), SSK (surface sample kit) slides, and MAS (microbial air sampler)
Petri dishes. T+5d analysis is not required. If Mike's analysis showed
colony growth above specified limits, he was to take digital documentary
images.]

Kaleri conducted his weekly IMS (inventory management system) tagups with
ground specialists, discussing open issues concerning identification of
equipment and storage locations for updating the IMS database. [TsUP has
now incorporated the recently completed EVA gloves audit in the IMS.]

Mike completed the regular task of transferring data files from the physical
exercise equipment to the MEC (medical equipment computer) via memory card
and RED log entries, for downlink on OCA comm. Later, he performed the
periodic transfer of accumulated data files from the wristband HRM (heart
rate monitor) receiver stations to the MEC for downlink, then deleting them
on the HRM.

The crew had their regular workouts on TVIS, RED exerciser and VELO cycle
with load trainer.

Sasha completed his regular weekly inspection of the BRPK air/condensate
water separator system. He also conducted the routine technical maintenance
of the SM's SOZh environment control & life support system, incl. the toilet
equipment (ASU). The regular daily preparation of the IMS "delta" file for
automatic export/import to update the database has now become part of the
Russian crew-discretionary task list.

Also added to the Russian task list was the incorporation of changes in the
Russian segment (RS) EMER-1 emergency procedures book (3 copies), to account
for the PEV (pressure equalization valve, Russian: KVD) at the FGB nadir
hatch (where Soyuz 8S will dock) being switched to "closed".

Mike Foale had the weekly crew conference with the JSC Astronaut Office and
the Increment 9 crew.

The crew conducted their weekly tag-up with the ISS Flight Director, via
S-band.

Mike Foale's successful maneuvering of the Canadian SSRMS (Space Station
Remote Manipulator System) yesterday allowed the ground to perform the
scheduled overnight ISS external/portside camera survey. [The survey
included the FGB pressurized compartment ("Ball")/PMA interface, Node/PMA-3
interface, Z1/P6 truss interface, Aft P6 radiator, P6 Ku antenna dish, US
Lab, S0 truss, MBS (mobile service system), P1 Nadir H-fixture, and the P1
radiator. Preliminary findings: a black mark on the back of the Ku-band
dish, a possible tear in the MLI (multi-layer insulation) blanket on the
SSRMS, a possible MLI problem on the P1 truss, and a possible obstruction on
an SPDM-H (Special Purpose Dexterous Manipulator) handle. Further in-depth
analysis and comparison of the imagery is underway.]

MCC-Houston has agreed to extend Russia's loan of the U.S. laptop #6075,
formerly SSC6 (station support computer #6), until the delivery of a new
laptop on Progress 14P. [The IBM ThinkPad 760XD was borrowed for the MBI-8
Profilaktika ("countermeasures") fitness assessment and may now be used in
lieu of laptop 3, currently operating with the SM Central Post computer #1
(KTsP1) instead of the failed Russian laptop 1.]

Today's CEO (Crew Earth Observations) targets, limited in XPOP attitude by
flight rule constraints on the use of the science window, which is available
for only ~1/4 of each orbit when not facing forward (in "ram"), were Smoke
pall, Thailand (looking right for obliques of an extensive smoke pall
drifting south into the Gulf of Siam. Interestingly, it is known that
Chinese naval and mercantile fleets, during the 7th, 12th and 15th centuries
at least, relied on messages from a network of carrier pigeons flying in
from S Asia. News of large, persistent smoke palls [over Burma, Malaya,
Sumatra, Kalimantan, Java] indicated mean wind conditions in the early
spring in this region far to the south of Beijing. Knowledge of spring wind
conditions in turn helped to optimize the timing of fleet sailings. This
method of returning meteorological information to Beijing persisted until
about 1440-1450 A.D.), Internal waves, Patagonia (looking slightly left of
track towards the early morning glint point for images of internal waves in
the wide swath of shallow seas between Buenos Aires and the Falkland
Islands), and Soufriere Volcano, Caribbean (this volcano continues to erupt.
Looking left to see Montserrat island about two degrees off track. The long
smoke and ash plume stretches west towards Puerto Rico).

CEO images can be viewed at the websites.

http://eol.jsc.nasa.gov
http://earthobservatory.nasa.gov

See also the website "Space Station Challenge" at
http://voyager.cet.edu/iss/

U.S. and Russian Segment Status (as of today, 2:20pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

Elektron O2 generator is Off. Vozdukh CO2 scrubber is On. U.S. CDRA CO2
scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant
control subsystem) is operating. SM Gas Analyzer has been calibrated and is
used for ppO2 and ppCO2 monitoring. MCA (major constituents analyzer) is in
Life Extending Mode (LEM). BMP Harmful Impurities unit: absorbent bed #1 in
Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On, SKV-2
is Off (repair now completed; to be tested ASAP).


SM Working Compartment: Pressure (mmHg) -- 742; temperature (deg C) --
26.0; ppO2 (mmHg) -- 139.7; ppCO2 (mmHg) -- 3.8;
SM Transfer Compartment: Pressure (mmHg) -- 747; temperature (deg C) --
20.3.
FGB Cabin: Pressure (mmHg) -- 744; temperature (deg C) -- 23.3.
Node: Pressure (mmHg) -- 740.7; temperature (deg C) -- 24.7 (shell); ppO2
(mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
U.S. Lab: Pressure (mmHg) -- 742.29; temperature (deg C) -- 24.7; ppO2
(mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
Joint Airlock (Equip. Lock): Pressure (mmHg) -- 744.3; temperature (deg
C) -- 27.7; shell heater temp (deg C) -- 25.7, ppO2 (mmHg) -- n/a; ppCO2
(mmHg) -- n/a.


PMA-1: Shell heater temp (deg C) -- 23.6
PMA-2: Shell heater temp (deg C) -- 9.6

(n/a = data not available)

Electrical Power Systems (EPS):

Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B
both in Autotrack (suntracking) and bias-angled 43 deg. for drag reduction
("sun slicer")
SM batteries: Battery #8 is off-line; all other batteries (7) are in
"Partial Charge" mode.
FGB batteries: Battery #6 is off-line (capacity restoration mode, ROM); all
other batteries (5) are in "Partial Charge" mode.
Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
INT-2 is operating; INT-1 is Off.
EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is Off; PL-2 MDM is Operational.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1
dropped out 11/22/03).
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3
dropped out 10/22/03).

Propulsion System:

Total propellant load available: 3948 kg (8704 lb) as of 3/5/04 [SM(755) +
FGB(2534) + Progress M(0) + Progress M-1(659)]. (Capability: SM -- 860 kg;
FGB -- 6120 kg).

Attitude Control Systems:

3 CMGs on-line (CMG-1 failed).
State vector source -- U.S. SIGI-1 (GPS)
Attitude source -- U.S. SIGI-1 (GPS)
Angular rate source -- RGA-1

Flight Attitude:

XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: 0.5 deg,
pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist), until 3/28.

Communications & Tracking Systems:

FGB MDM-1 is powered Off; FGB MDM-2 is operational.
All other Russian communications & tracking systems are nominal.
*S-band is operating nominally (on string 2).
Ku-band is operating nominally.
Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
Video subsystem operating nominally.
HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

SSRMS/Canadarm2 based at Lab PDGF/LEE A, powered on both strings.
MBS: KA power on both strings.
MT: latched and mated at WS4.
POA: KA power on both strings.
RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is
Off.

ISS Orbit (as of this morning, 7:23am EST [= epoch]):

Mean altitude -- 367.5 km
Apogee -- 374.7km
Perigee -- 360.4 km
Period -- 91.90 min.
Inclination (to Equator) -- 51.6288 deg
Eccentricity -- 0.0010563
Solar Beta Angle -- ~47 deg (magnitude decreasing)
Orbits per 24-hr. day -- 15.67
Mean altitude loss last 24 hours -- 100 m
Revolutions since FGB/Zarya launch (Nov. '98) -- 30424

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times,
see http://www.hq.nasa.gov/osf/station/viewing/issvis.html



--
---------------------------

Jacques :-)

www.spacepatches.info




 




Thread Tools
Display Modes

Posting Rules
You may not post new threads
You may not post replies
You may not post attachments
You may not edit your posts

vB code is On
Smilies are On
[IMG] code is On
HTML code is Off
Forum Jump

Similar Threads
Thread Thread Starter Forum Replies Last Post
ISS Status Report No. 13 - 2004 Jacques van Oene Space Station 0 March 13th 04 10:46 AM
ISS Status Report No. 06 - 2004 Jacques van Oene Space Station 0 January 29th 04 05:50 PM
ISS Status Report No. 04 - 2004 Jacques van Oene Space Station 0 January 17th 04 01:01 PM
ISS Status Report No. 01 - 2004 Jacques van Oene Space Station 0 January 4th 04 02:43 PM
Ed Lu Letter from Space #6 Jacques van Oene Space Station 0 July 4th 03 11:10 AM


All times are GMT +1. The time now is 05:25 AM.


Powered by vBulletin® Version 3.6.4
Copyright ©2000 - 2025, Jelsoft Enterprises Ltd.
Copyright ©2004-2025 SpaceBanter.com.
The comments are property of their posters.